show that p(-5/7,-2sqrt(6)/7) is on the unit circle
Answer by Alan3354(21605)

show that p(-5/7,-2sqrt(6)/7) is on the unit circle
r^2 = (5/7)^2 + (2sqrt(6)/7)^2
= 25/49 + 4*6/49
= 49/49
= 1
radius = 1, so it's on the circle.